Mr. Samson can sort the bins in 12 days. If Joe Helps him, it takes only 4 days. How long would it take Joe to sort the bins himself
Answer by stanbon(75887)

Mr. Samson can sort the bins in 12 days. If Joe Helps him, it takes only 4 days. How long would it take Joe to sort the bins himself
-----------------
Samson DATA:
Time = 12 day/job : Rate = 1/12 job/day
-----------------------
Together DATA:
Time = 4 days/job ; Rate = 1/4 job/day
--------------------
Joe DATA:
Time = x days/job ; Rate = 1/x job/day
----------------------
EQUATION:
rate + rate = together rate
1/12 + 1/x = 1/4
Multiply thru by 12x to get:
x + 12 = 3x
2x = 12
x = 6 (time Joe would need to do the job alone)
=================
Cheers,
Stan H.